# A Deep Neural Network Approach to Compact Source Removal

### v1

@@ 1 Introduction
In this paper, we present a deep-learning tool, designed to remove point sources from the photometric observations of the PACS instrument (Poglitsch et al., 2010) on-board of the Herschel Space Observatory (Pilbratt et al., 2010). This work is part of our Novel Evolutionary Model for the Early Stages of stars with Intelligent Systems (NEMESIS [^1] https://nemesis.univie.ac.at) project, and a continuation and extension of the new Herschel/PACS catalog (Marton et al., 2024). The tool can be trained for the photometric data of other instruments on other spacecrafts, as well, thanks to the modified U-Net architecture with partial convolutional layers integrated within an inpainting framework, central to our approach. In Section 2, we describe the data used for the different steps of preprocessing, training and validation. In Section 3, we give a detailed description of the deep-learning algorithm, the architecture, the implementation and the photometric methods used for validation. In Section 4, we present the performance of the method, and in Section 5, we discuss the photometric accuracy and the image properties of the resulting data after applying our method to various examples.
@@ 2.2 Point Spread Function
For 60′′s-1 scan speeds, the shape of the PSF is more elongated. However, with different scan angles, the elongation occurs in different directions, which can be handled using circular masks, as discussed in Section 2.4.4.
@@ 2.3 Simulations
As mentioned earlier, with the two Observing Modes and the three bands, we have a total of six different types of data to process. However, since the three bands within the modes are more similar in appearance in the image cutouts, we decided to merge the data for each mode. This way, we only need to train two models, one for the Scan Map mode and one for the Parallel mode, and allow the network to learn the differences between the bands using masks designed specifically for the bands, as described in Section 2.4.4. The need to handle Scan Map and Parallel modes separately arises from their resolutional differences.
@@ 2.3.1 Data Generation Process
    A corresponding .csv file was also created, which included the equatorial coordinates of right ascension (Ra) and declination (Dec) in the J2000.0 reference frame of the injected sources, with the flux density levels used in mJy.
@@ 2.3.2 Dataset Partitioning
Next, we followed the steps described in Section 2.3.1 to create the simulations for each selected obsid field. We then divided the list of obsids into training, validation and test sets, ensuring that each of the three datasets contained at least one field from each of the custom categories. The training, validation and test sets are disjoint, since we used different obsids in each set. Table 2 shows the distribution of sources and fields in the Scan Map and Parallel datasets, and additional details about the fields used can be found in Table 6, Table 7 and Table 8 in the Appendix. The reason why we did not use simulations from the Parallel Red dataset for training is because it has the same PSF as the Scanmap version, thus only test simulations were created.
@@ 3 Methods
We developed a neural network architecture grounded in an inpainting approach to address the challenge of compact source removal and extraction. Inpainting refers to a technique used to reconstruct missing or masked parts of an image based on the surrounding information. For our method, we use dynamically constructed masks with the SNR values, as described in Section 2.4.4.
@@ 3.4 Astropy and Aperture Photometry
We also performed aperture photometry correction to account for the portion of the PSF that lies outside the aperture. This correction is important for obtaining a more accurate flux measurement, as parts of the PSF of the sources fall outside the defined aperture radius.
@@ 3.5.1 Setup and Hyperparameters
The training inputs were 1-channel images of size $96\times 96$ pixels, as detailed in Section 2.4.5. The model was trained using a batch size of 64.
@@ 3.5.2 Epochs and Validation
The training of our model followed a systematic approach, after each epoch of training on the training dataset, a validation loop was conducted on the separate validation dataset. During these validation loops, we closely observed the performance of the model using the final loss function (Equation 7).
@@ 4 Results
Our main result is the open-source tool we developed and made available for the community to use in removing compact sources, hosted at our GitHub repository  [^5] https://github.com/xMediMadix/compact-source-removal. Users can employ the pre-trained model for source removal from Herschel images, or re-train the model on custom datasets. The repository includes three main configuration files, each designed for one of the tasks described below, along with a detailed README.md file explaining how to utilize the code effectively. In Section 5, we will evaluate our model’s performance across various aspects. This includes an accuracy comparison of the photometry pipeline we used on standard star models, a comparison between the measured flux densities of extracted sources and the known values of the injected PSFs, and an analysis of how the model impacts the power spectrum and the pixel intensity distributions of the maps.
@@ 5.1 Comparison with Standard Star Models
Studies by Balog et al. (2014) and Klaas et al. (2018) have previously examined the flux calibration of the PACS photometer. We applied our network to estimate the background and extract the sources, then performed aperture photometry as discussed in Section 3.4 to measure the flux density of selected standard stars. We used the fluxes provided in these studies as references to evaluate both the accuracy of the aperture_photometry function and the source removal performance of our network. Table 5 summarizes the comparison of the reference flux densities from the studies with the measured flux densities of the sources extracted by our network across the $70\mu\mathrm{m}$, $100\mu\mathrm{m}$, and $160\mu\mathrm{m}$ bands.
@@ 5.2 Photometric Accuracy
As described in Section 2.3.2, we preserved several fields for testing purposes, with the numbers shown in Table 2. These fields were not used in either the training or validation loops, making them suitable for testing our model on unseen cases.
@@
We applied the same processing steps to our test dataset as we did to the other datasets: each image was per-image normalized, mask sizes were determined using the same parameters, and all inputs were passed through our pre-trained model to create source-free, background-only images for all injected sources. We obtained the extracted source by subtracting the model’s predicted background from the original input image. Next, we performed aperture photometry on this background-subtracted source with various aperture sizes, ranging from 4 to 18 arcseconds, to measure the extracted flux density. We also performed circular annulus background estimation on the input images, as described in Section 3.4, and compared these flux density values to the known injected flux densities by calculating their ratios.
@@ 5.3 Image comparison
Figure 13: Left column: Direction independent Fourier transform power spectra of 3 different fields before adding artificial sources (blue line), with the artificial sources (orange) and after their removal (green). Power spectra are normalized. Middle column: Difference of the power spectrum before adding sources and after their removal. Right column: Probability density functions before adding artificial sources and after their removal. Blue histograms show the PDF of the unmodified map, red histograms show the PDFs of the predicted maps.
@@ 5.3.1 Power spectrum
The Fourier transform power spectra of the different maps provide useful insight into the properties of the signal introduced by our tool. In Fig. 13 we demonstrate a comparison with one case in each filter band. These fields were chosen from the test dataset, and for each field, we generated random positions for the sources. For each random position, a random flux density and orientation were also assigned, ensuring that the fields represent the most diverse and variable cases possible. These maps are similar to the example shown in Fig. 9.
@@ Acknowledgements.
This work has received funding from the European Union’s Horizon 2020 research and innovation program under grant number 101004141. G.M. acknowledges support from the János Bolyai Research Scholarship of the Hungarian Academy of Sciences. This research was supported by the International Space Science Institute (ISSI) in Bern, through ISSI International Team project 521 selected in 2021, Revisiting Star Formation in the Era of Big Data (https://teams.issibern.ch/starformation/).

### v2

@@ 1 Introduction
In this paper, we present a deep-learning tool, designed to remove point sources from the photometric observations of the PACS instrument (Poglitsch et al., 2010) on-board of the Herschel Space Observatory (Pilbratt et al., 2010). This work is part of our Novel Evolutionary Model for the Early Stages of stars with Intelligent Systems (NEMESIS [^1] https://nemesis.univie.ac.at) project, and a continuation and extension of the new Herschel/PACS catalog (Marton et al., 2024). The tool can be trained for the photometric data of other instruments on other spacecrafts, as well, thanks to the modified U-Net (Ronneberger et al., 2015) architecture with partial convolutional layers integrated within an inpainting framework, central to our approach. Section 2 outlines the data used across the different stages of preprocessing, training, and validation. A detailed description of the deep-learning algorithm, including its architecture, implementation, and the photometric methods employed for validation, is provided in Sect. 3. The performance of the method is then presented in Sect. 4, followed by a discussion in Sect. 5 on photometric accuracy and the image properties of the resulting data after applying our approach to various examples.
@@ 2.2 Point Spread Function
For 60′′s-1 scan speeds, the shape of the PSF is more elongated. However, with different scan angles, the elongation occurs in different directions, which can be handled using circular masks, as discussed in Sect. 2.4.4.
@@ 2.3 Simulations
As mentioned earlier, with the two Observing Modes and the three bands, we have a total of six different types of data to process. However, since the three bands within the modes are more similar in appearance in the image cutouts, we decided to merge the data for each mode. This way, we only need to train two models, one for the Scan Map mode and one for the Parallel mode, and allow the network to learn the differences between the bands using masks designed specifically for the bands, as described in Sect. 2.4.4. The need to handle Scan Map and Parallel modes separately arises from their resolutional differences.
@@ 2.3.1 Data Generation Process
    A corresponding .csv file was also created, which included the equatorial coordinates of right ascension (R.A.) and declination (Dec) in the J2000.0 reference frame of the injected sources, with the flux density levels used in mJy.
@@ 2.3.2 Dataset Partitioning
Next, we followed the steps described in Sect. 2.3.1 to create the simulations for each selected obsid field. We then divided the list of obsids into training, validation and test sets, ensuring that each of the three datasets contained at least one field from each of the custom categories. The training, validation and test sets are disjoint, since we used different obsids in each set. Table 2 shows the distribution of sources and fields in the Scan Map and Parallel datasets, and additional details about the fields used can be found in Table 6, Table 7 and Table 8 in the Appendix. The reason why we did not use simulations from the Parallel Red dataset for training is because it has the same PSF as the Scanmap version, thus only test simulations were created.
@@ 3 Methods
We developed a neural network architecture grounded in an inpainting approach to address the challenge of compact source removal and extraction. Inpainting refers to a technique used to reconstruct missing or masked parts of an image based on the surrounding information. For our method, we use dynamically constructed masks with the SNR values, as described in Sect. 2.4.4.
@@ 3.4 Astropy and Aperture Photometry
We also performed aperture photometry correction to account for the portion of the PSF that lies outside the aperture. This correction is important for obtaining a more accurate flux measurement, as parts of the PSF of the sources fall outside the defined aperture radius. Additionally, the shape of the PSF depends on the source’s spectral energy distribution (SED). In our simulations, we used the PSF based on Vesta, which has an effective temperature that differs significantly from that of typical standard stars. To address this, we applied aperture corrections specific to Vesta for our simulated sources, while using different aperture corrections for comparison to standard star models in Sect. 5.1 (see Lutz (2015)).
@@ 3.5.1 Setup and Hyperparameters
The training inputs were 1-channel images of size $96\times 96$ pixels, as detailed in Sect. 2.4.5. The model was trained using a batch size of 64.
@@ 3.5.2 Epochs and Validation
The training of our model followed a systematic approach, after each epoch of training on the training dataset, a validation loop was conducted on the separate validation dataset. During these validation loops, we closely observed the performance of the model using the final loss function Eq. 7.
@@ 4 Results
Our main result is the open-source tool we developed and made available for the community to use in removing compact sources, hosted at our GitHub repository  [^5] https://github.com/xMediMadix/compact-source-removal. Users can employ the pre-trained model for source removal from Herschel images, or re-train the model on custom datasets. The repository includes three main configuration files, each designed for one of the tasks described below, along with a detailed README.md file explaining how to utilize the code effectively. In Sect. 5, we will evaluate our model’s performance across various aspects. This includes an accuracy comparison of the photometry pipeline we used on standard star models, a comparison between the measured flux densities of extracted sources and the known values of the injected PSFs, and an analysis of how the model impacts the power spectrum and the pixel intensity distributions of the maps.
@@ 5.1 Comparison with Standard Star Models
Studies by Balog et al. (2014) and Klaas et al. (2018) have previously examined the flux calibration of the PACS photometer. We applied our network to estimate the background and extract the sources, then performed aperture photometry as discussed in Sect. 3.4 to measure the flux density of selected standard stars. We used the fluxes provided in these studies as references to evaluate both the accuracy of the aperture_photometry function and the source removal performance of our network. Table 5 summarizes the comparison of the reference flux densities from the studies with the measured flux densities of the sources extracted by our network across the $70\mu\mathrm{m}$, $100\mu\mathrm{m}$, and $160\mu\mathrm{m}$ bands.
@@ 5.2 Photometric Accuracy
As described in Sect. 2.3.2, we preserved several fields for testing purposes, with the numbers shown in Table 2. These fields were not used in either the training or validation loops, making them suitable for testing our model on unseen cases.
@@
We applied the same processing steps to our test dataset as we did to the other datasets: each image was per-image normalized, mask sizes were determined using the same parameters, and all inputs were passed through our pre-trained model to create source-free, background-only images for all injected sources. We obtained the extracted source by subtracting the model’s predicted background from the original input image. Next, we performed aperture photometry on this background-subtracted source with various aperture sizes, ranging from 4 to 18 arcseconds, to measure the extracted flux density. We also performed circular annulus background estimation on the input images, as described in Sect. 3.4, and compared these flux density values to the known injected flux densities by calculating their ratios.
@@ 5.3 Image comparison
Figure 13: From top to bottom, the three rows correspond to the Scan Map Blue, Green, and Red fields, with observation IDs 1342216409, 1342244224, 1342249136, respectively. Left column: Direction independent Fourier transform power spectra of each field before adding artificial sources (blue line), with the artificial sources (orange) and after their removal (green). Power spectra are normalized. Middle column: Difference of the power spectrum before adding sources and after their removal for each field. Right column: Probability density functions before adding artificial sources (blue histograms) and after their removal (red histograms).
@@ 5.3.1 Power spectrum
The Fourier transform power spectra of the different maps provide useful insight into the properties of the signal introduced by our tool. In Fig. 13 we demonstrate a comparison with one case in each filter band in the Scan Map dataset. These fields were chosen from the test dataset, and for each field, we generated random positions for the sources. For each random position, a random flux density and orientation were also assigned, ensuring that the fields represent the most diverse and variable cases possible. These maps are similar to the example shown in Fig. 9.
@@ Acknowledgements.
This work has received funding from the European Union’s Horizon 2020 research and innovation program under grant number 101004141. G.M. acknowledges support from the János Bolyai Research Scholarship of the Hungarian Academy of Sciences. This research was supported by the International Space Science Institute (ISSI) in Bern, through ISSI International Team project 521 selected in 2021, Revisiting Star Formation in the Era of Big Data (https://teams.issibern.ch/starformation/). We are grateful to our referee for the insightful comments and suggestions that have greatly contributed to further improving this paper.